How much 151 Rum, which is 75.5% alcohol needs to be mixed with Rum 80, which is 40% alcohol to obtain a 1000ml 65% alcohol solution?
Answer by stanbon(48558)

How much 151 Rum, which is 75.5% alcohol needs to be mixed with Rum 80, which is 40% alcohol to obtain a 1000ml 65% alcohol solution?
----------------------------------------------------
Equation:
alcohol + alcohol = alcohol
0.755x + 0.40(1000-x) = 0.65(1000)
Multiply thru by 1000
755x + 400*1000-400x = 650*1000
355x = 250000
x = 704.23 ml (amt. of 75.5% alcohol in the mixture)
----------------
1000-x = 285.77 ml (amt. of 40% alcohol in the mixture)
===========================================================
Cheers,
Stan H.